# Knowledge Distillation on Spatial-Temporal Graph Convolutional Network for Traffic Prediction

### v2

@@ 4.5 Pruning and Fine-Tuning Analysis
In this section, the focus is on modifying the student network through a pruning algorithm, intentionally applying the same algorithm to the teacher network. The objective is to create a student network with parameters comparable to the one used in knowledge distillation experiments. The goal is to showcase that a consciously pruned student network, guided by Algorithm 1, exhibits superior learning capabilities compared to a predetermined student network used in knowledge distillation. The experiment’s outcomes are summarized in Table 7 and Figures 4 and 5.
@@
Figure 4: Results illustrating the impact of employing the pruning algorithm to enhance knowledge distillation. The left chart in the top row corresponds to the teacher, while the right chart in the top row pertains to a student without knowledge distillation. In the second row, the left chart represents a student with knowledge distillation but without pruning, and the right chart represents a student with both knowledge distillation and pruning.
[…]
(a) PeMSD7
(b) PeMSD8
@@
Next, we conduct a comparative analysis between the presented pruning algorithm and a non-distillation scenario from [6]. Previous research has confirmed that incorporating knowledge distillation enhances accuracy while utilizing a more straightforward model. Building upon this, our current study extends the concept, illustrating that Algorithm 1 yields a student architecture that performs better during training with KD. We compare Algorithm 1 with [6] using a base pruning model from Table 1. The results depicted in Table 8 underscore that our pruning algorithm consistently outperforms the alternative across various pruning percentages.
Table 8: Comparing the performance of Algorithm 1 with the algorithm proposed in [6].
[…]
In the PeMSD7 dataset, our proposed pruned W $L_{\text{STCD}}$ (pruned with $L_{\text{STCD}}$) consistently exhibits superior performance compared to the pruned W/O KD (without knowledge distillation) approach. This superiority is evident in metrics such as mean absolute percentage error (MAPE), mean absolute error (MAE), and root mean square error (RMSE) across different time intervals (15 min, 30 min, and 45 min). For example, at a pruning level of $25\%$, pruned W $L_{\text{STCD}}$ achieves lower MAPE values (5.950, 8.654, 10.906) compared to pruned W/O KD (6.185, 8.981, 11.060) for 15 min, 30 min, and 45 min intervals, respectively. Similar trends are observed at higher pruning levels of $50\%$ and $75\%$, where pruned W $L_{\text{STCD}}$ consistently outperforms pruned W/O KD across all evaluation metrics. The superiority of our algorithm is further emphasized in the PeMSD8 dataset. At various pruning levels, pruned W $L_{\text{STCD}}$ consistently records lower MAPE, MAE, and RMSE values compared to pruned W/O KD. Remarkably, even at aggressive pruning levels of $75\%$, pruned W $L_{\text{STCD}}$ maintains competitive performance, showcasing the effectiveness and robustness of our proposed pruning algorithm. Overall, the findings affirm that our introduced pruning algorithm plays a pivotal role in enhancing learning through knowledge distillation, significantly improving neural network accuracy by employing fewer parameters and simpler structures. In the executions related to comparing our pruning algorithm with the one presented in [6], pruning operations are performed after every two cycles, and the final network cycle undergoes fine-tuning (see Table 8).
4.5.1 Hyperparameters
The hyperparameters for the proposed pruning algorithm are presented in Table 9. The table categorizes hyperparameters based on the percentage of pruning applied, as well as batch size, learning rate, and three alpha values ($\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$), specific to the PeMSD7 and PeMSD8 datasets. For instance, in the PeMSD7 row with a pruning percentage of $97\%$, only $3\%$ of the model parameters are retained. The associated hyperparameters include a batch size of 25, a learning rate of $1\times 10^{-3}$, and alpha values $\alpha_{1}=0.099$, $\alpha_{2}=0.091$, and $\alpha_{3}=0.531$. The alpha values play distinct roles in the loss function equations. For $\alpha_{1}$, it defines the conditions under which different loss terms are applied in $L_{\text{ORD}}$ as specified in Equation 6. Meanwhile, $\alpha_{2}$ and $\alpha_{3}$ are used as weights in Equation 11, influencing the contributions of $L_{\text{SCD}}$ and $L_{\text{TCD}}$ relative to $L_{\text{ORD}}$.
Table 9: Hyperparameters of the proposed pruning algorithm
@@ 5.2 Future Works
As we progress, we have the opportunity to explore our cost function $L_{\text{STCD}}$ and Algorithm 1 on alternative spatiotemporal graph neural networks, such as ASTGCN. Furthermore, considering the method proposed in [32] and [26], we can explore whether using multiple teachers in Algorithm 1 and knowledge distillation with $L_{\text{STCD}}$ can lead to a model with higher accuracy. Alternatively, we can develop a model capable of executing the tasks performed by all its teachers with increased accuracy and reduced execution time compared to the model in [26]. It is possible to enhance the concept of knowledge distillation based on the cost function $L_{\text{STCD}}$ by incorporating relationship-based approaches. This improvement allows for the consideration of student and teacher networks with varying numbers of layers.

### v4

@@ 4.1 Hyperparameters
Table 3: Hyperparameters of the proposed pruning algorithm
[…]
4.2 Knowledge Distillation Analysis
[…]
We conducted a comparative analysis of our final loss function, $L_{\text{STCD}}$, against several benchmarks: the baseline student model without knowledge distillation (KD), ARIMA, Historical Average, ASTGCNGuo et al. [2019], Li et al. [2021], and other established approaches Bian et al. [2021], Huang and Wang [2017], Zagoruyko and Komodakis [2016], Romero et al. [2014], Yang et al. [2020]. The results are summarized in Table 4.
In reviewing the table, it is evident that for the PeMSD7 dataset, $L_{\text{STCD}}$ consistently achieves lower Mean Absolute Error (MAE) and Root Mean Square Error (RMSE) for both 15-minute and 30-minute predictions. Similarly, on the PeMSD8 dataset, $L_{\text{STCD}}$ demonstrates superior performance across various metrics when compared to $L_{\text{RD(L2)}}$, $L_{\text{RD(KL)}}$, and $L_{\text{SKD}}$, as well as to lightweight models like ASTGCN. Figure 3 depict hidden layer representations for 50 test data points, with each point representing an individual instance. Red dots represent the student, while blue dots denote either the teacher or the baseline pruning network (utilized for specific pruning results). Proximity or overlap between red and blue dots serves as an indicator of the model’s success. Ideally, effective methods show red and blue dots close to or overlapping each other. Figure 3 illustrates better convergence and pattern imitation of the $L_{\text{STCD}}$ approach compared to other approaches in both knowledge bases.
Table 4: Comparison of our approach $L_{\text{STCD}}$ with Historical Average, ARIMA, $L_{\text{RD(L2)}}$, $L_{\text{RD(KL)}}$ and $L_{\text{SKD}}$ on PeMSD7 and PeMSD8
Models PeMSD7 MAPE MAE RMSE 15 min 30 min 45 min 15 min 30 min 45 min 15 min 30 min 45 min Teacher 5.223 7.316 8.739 2.230 3.010 3.565 4.097 5.752 6.834 HA 10.610 4.010 7.200 ARIMA 13.400 14.010 15.010 5.570 5.940 6.270 9.000 9.220 9.430 ASTGCN 7.250 8.670 9.730 2.850 3.350 3.700 5.150 6.120 6.770 Student without KD 6.423 9.685 12.298 2.666 3.868 4.799 4.649 6.938 8.610 Student $L_{\text{RD(L2)}}$ 6.379 9.661 12.474 2.768 4.214 5.527 4.709 7.185 9.178 Student $L_{\text{RD(KL)}}$ 6.411 9.527 11.894 2.700 3.938 4.918 4.672 6.984 8.657 Student $L_{\text{SKD}}$ 6.282 9.503 11.902 2.762 4.092 5.101 4.735 7.161 8.923 Student $L_{\text{STCD}}$(ours) 6.078 9.043 11.488 2.615 3.776 4.754 4.537 6.678 8.344
Models PeMSD8 MAPE MAE RMSE 15 min 30 min 45 min 15 min 30 min 45 min 15 min 30 min 45 min Teacher 2.293 3.239 3.925 1.211 1.665 2.031 2.524 3.501 4.081 HA 3.940 1.980 4.110 ARIMA 5.110 5.210 5.460 1.900 2.120 2.430 4.870 5.240 5.630 ASTGCN 3.160 3.590 3.980 1.490 1.670 1.81 3.180 3.690 3.920 Student without KD 2.967 4.035 4.734 1.472 1.956 2.294 2.988 4.090 4.706 Student $L_{\text{RD(L2)}}$ 2.812 3.976 4.908 1.426 1.953 2.387 2.839 3.978 4.733 Student $L_{\text{RD(KL)}}$ 2.964 4.179 5.057 1.507 2.125 2.575 2.895 3.971 4.664 Student $L_{\text{SKD}}$ 2.726 3.819 4.553 1.409 1.944 2.317 2.909 4.139 4.894 Student $L_{\text{STCD}}$(ours) 2.491 3.375 4.067 1.281 1.719 2.052 2.716 3.707 4.385
(a) PeMSD7
(b) PeMSD8
Figure 3: Comparison of our spatial-temporal correlation distillation loss function $L_{\text{STCD}}$ with the $L_{\text{RD(L2)}}$, $L_{\text{RD(KL)}}$ and $L_{\text{SKD}}$ loss functions. Subfigure 3(a) shows results for PeMSD7 dataset, and Subfigure 3(b) shows for PeMSD8. The left chart in the top row corresponds to the teacher, while the middle chart in the top row pertains to a student without knowledge distillation. In the second row, the right chart represents our loss function, and the other charts indicate different loss functions.
We also conducted an ablation study on our final loss function, $L_{\text{STCD}}$; response-based distillation, $L_{\text{ORD}}$; and other variants, including $L_{\text{TCD}}$, $L_{\text{SCD}}$, along with the baseline Student without KD (student trained without KD), and the findings are outlined in Table 5. Additionally, we report the ratio of attention to teacher output instead of the dataset target in Table 6.
Table 5: Ablation study for our $L_{\text{STCD}}$ loss function on PeMSD7 and PeMSD8
Models PeMSD7 MAPE MAE RMSE 15 min 30 min 45 min 15 min 30 min 45 min 15 min 30 min 45 min Teacher 5.223 7.316 8.739 2.230 3.010 3.565 4.097 5.752 6.834 Student without KD 6.423 9.685 12.298 2.666 3.868 4.799 4.649 6.938 8.610 Student $L_{\text{ORD}}$ 6.411 9.516 12.104 2.762 4.091 5.221 4.645 6.847 8.569 Student $L_{\text{TCD}}$ 6.320 9.411 11.667 2.730 3.966 4.853 4.678 6.899 8.512 Student $L_{\text{SCD}}$ 6.326 9.193 11.380 2.743 3.957 4.866 4.645 6.853 8.476 Student $L_{\text{STCD}}$ 6.078 9.043 11.488 2.615 3.776 4.754 4.537 6.678 8.344
Models PeMSD8 MAPE MAE RMSE 15 min 30 min 45 min 15 min 30 min 45 min 15 min 30 min 45 min Teacher 2.293 3.239 3.925 1.211 1.665 2.031 2.524 3.501 4.081 Student without KD 2.967 4.035 4.734 1.472 1.956 2.294 2.988 4.090 4.706 Student $L_{\text{ORD}}$ 2.661 3.717 4.553 1.363 1.885 2.285 2.788 3.862 4.573 Student $L_{\text{TCD}}$ 2.509 3.497 4.215 1.296 1.747 2.063 2.665 3.716 4.406 Student $L_{\text{SCD}}$ 2.619 3.680 4.457 1.355 1.850 2.214 2.722 3.778 4.478 Student $L_{\text{STCD}}$ 2.491 3.375 4.067 1.281 1.719 2.052 2.716 3.707 4.385
@@ 4.3 Pruning and Fine-Tuning Analysis
[…]
@@
Previous research has confirmed that incorporating knowledge distillation enhances accuracy while utilizing a more straightforward model. Building upon this, our current study extends the concept, illustrating that Algorithm 1 yields a student architecture that performs better during training with KD. We compare Algorithm 1 with Molchanov et al. [2019] using a base pruning model from Table 1 (second row). The results depicted in Table 8 underscore that our pruning algorithm consistently outperforms the alternative across various pruning percentages. Overall, the findings affirm that our introduced pruning algorithm plays a pivotal role in enhancing learning through knowledge distillation, significantly improving neural network accuracy by employing fewer parameters and simpler structures.
Table 8: Comparison of the performance between Algorithm 1 and the pruning algorithm that does not utilize knowledge distillation.
@@ 5.2 Future Works
As we progress, we have the opportunity to explore our cost function $L_{\text{STCD}}$ and Algorithm 1 on alternative spatiotemporal graph neural networks, such as ASTGCN. Furthermore, considering the method proposed in Luo et al. [2019] and Jing et al. [2021], we can explore whether using multiple teachers in Algorithm 1 and knowledge distillation with $L_{\text{STCD}}$ can lead to a model with higher accuracy. Alternatively, we can develop a model capable of executing the tasks performed by all its teachers with increased accuracy and reduced execution time compared to the model in Jing et al. [2021]. It is possible to enhance the concept of knowledge distillation based on the cost function $L_{\text{STCD}}$ by incorporating relationship-based approaches. This improvement allows for the consideration of student and teacher networks with varying numbers of layers.
Additionally, improving the interpretability of distilled models can significantly increase trust and provide deeper insights into traffic patterns. Understanding how the student network makes its predictions can help identify potential improvements and uncover new traffic dynamics. For example, employing techniques like the ones proposed in Chen et al. [2023], which focus on enhancing the transparency of deep learning models, could be beneficial for our traffic prediction models.
Finally, developing advanced pruning techniques that consider structural properties of the network can lead to more efficient and compact models without sacrificing accuracy. By focusing on the importance of connections within the network, as seen in recent methods such as those described in Wang et al. [2023], we can achieve significant reductions in model complexity while maintaining high performance. This approach aligns with our goal of optimizing the student network for real-time traffic prediction on resource-constrained devices.
In summary, future work can expand on our current framework by integrating adaptive learning mechanisms, improving model interpretability, and utilizing advanced pruning techniques, all of which will contribute to creating more robust and efficient traffic prediction models.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Data availability
Data will be made available on request.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used ChatGPT 3.5 in order to refine language. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.